# Globally defined Carroll symmetry of gravitational waves

### v2

@@ IV Schwarzian aspects
For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative of a function $f$ of is some real variable $x$ Ovsienko ; Vlahakis ; ZZH ,
@@
cf. (II.3) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
@@
The form (II.7) of $S$ then implies that $P$ determines also the second solution we called $Q$.
@@
instead of $f=P/Q=1/S$ as in (IV.6)  [^4] The singularities of $S$ has will be discussed in the next section for PT. The new cast (IV.8) then yields, once again, a SL eqn. (IV.2) whose profile is now determined by the Souriau matrix  [^5] Eqn. (IV.10) can be viewed as a equation for the Souriau matrix $S$.,
@@
Somewhat surprisingly, inserting here $S$ from (II.7), the previous Sturm-Liouville equation (II.3) with the same profile $\mathcal{A}$} is recovered. These results indicate that the interchange
@@
of the Schwarzian. The difference between the two casts comes from the initial conditions. Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ depend also on the integral (II.7).
@@ V Illustration by the Pöschl - Teller profile
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (V.1)
The Sturm-Liouville equation (II.3) is solved by Legendre functions. DM trajectories arise when $m$ is an integer, and we get Legendre polynomials, $(-1)^{m}P_{m}(\tanh U)$ DM-1 which have $m$ zeros.
Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$, but diverges at the junction points. For $m=1$, for example,
@@
obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, is regular either in $I_{-}<0$ or in $I_{-}>0$, but diverges at $u=0$, as depicted in FIG.1a. Similarly for $m=2$,
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (V.3)
is singular where the denominator vanishes, as shown in FIG.1b.
Figure 1: For integer $m$ the Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories, written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
$Q_{m}=P_{m}S_{m}$ , a second solution (III.3) of the Sturm-Liouville equation (II.3) shown in FIG.1, is a Legendre function. It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits a VM behaviour where $P_{m}$ is DM.
The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (V.4)
The two SturmLiouville{Sturm-Liouville\;} solutions are shown in FIG. 1 in red and in green.
The $P(U)$ found above provide us with DM geodesics which sare consistent with (II.15). Their caustic behavior at the zeros of $P$ is manifest in FIG.2.
Figure 2: For a DM wave in $D=1$ transverse dimension Brinkmann trajectories focus to the caustic points determined by $P_{m}(U_{k})=0$, $k=1,\dots,m$. For $m=1$ there is one focal point at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
Coming to symmetries, the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
(a)                       (b)
Figure 3: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which is also a trajectory] is odd. (b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even. The only zero of ${\bf P}_{1}$ is at $U=0$.
(a)                       (b)
Figure 4: Infinitesmal Carroll generators (V) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd.
The globally defined Carroll symmetry generators (III.1) are
@@
VI Representation of the symmetry generators
The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory. In FIGs. 58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (V.5)-(V).
@@ VI Representation of the symmetry generators
Figure 7: Translations act along the $m=2$ DM trajectory by (V) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.7, the translation generator (with $c=1/2$) is, for $U=$-$\infty$ , $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 3 and 4 confirm that, consistently with (V.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
@@
Boosts along the $m=2$ trajectory are shown in. FIG.8. The longest vector on the left (in brown) is, consistently with (V.5) and (V) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$. Combining FIGs. 3 with FIG.5 provides us with the outfolding to 3dimensions, shown in FIG. 9.
Figure 9: The translations implemented along the trajectories with $m=1$ and $m=2$.
@@
Figure 10: Carroll boosts implemented along the trajectories with $m=1$ and $m=2$.
Carroll boosts implemented along the trajectories behave at pure translations at P(U)=0P(U)~=~0. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
@@ VII Conclusion
Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics. The general expression (III.5) which involves both DM and VM is one of our main results.
When the parameter $m$ is an integer, then the wavezone accommodates $m$ half-waves and we get DM. For $m=2\ell+1$ $P(U)$ is odd and we do get net displacement. For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.15) and confirmed by FIGs 2, 3a and 4a.
DM trajectories are distinguished by having zero momentum and can be found by putting ${\bm{p}}_{0}=0$ into (III.5). Turning on the momentum, ${\bm{p}}_{0}\neq 0$, distorts the trajectory to VM.
The merit of BJR coordinates is their simplicity Sou73 ; Carroll4GW . For our initial conditions (II.16), the “motion” reduces to a fixed point, given by the conserved boost momentum ${\bm{x}}_{0}={\bf k}$ in (II.10). The entire dynamics is carried by $P(u)$. The price to pay is that the BJR description works only between zeros of the $P$-matrix, and must then be fitted together. These same points, distinguished by the vanishing of $\det(P)$.
The advantage of the Brinkmann form (III.1) is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$ – another one of our principal results,
Examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to the whole spacetime. The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system. The Brinkmann trajectories are focused at the zeros of $\det(P)$, as shown in FIG.2.
The Carroll symmetry generators, first identified by Souriau Sou73 , look more complicated in Brinkmann coordinates as their BJR counterparts do. They are however defined globally. They are also directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by making use of the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5). The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation EZHRev ; GenRelGrav .
@@
Further insight is provided by the Schwarzian framework Ovsienko ; ZZH ; Vlahakis . This approach confirms that $Q=PS$ is indeed a global, independent, 2nd solution of the Sturm-Liouville problem which does not involve the locally defined Souriau matrix $S$.
The supersymmetric extension MSUSY sheds further light on the Memory Effect.
Our results go beyond $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which yields an analytic approximation of flyby ZelPol ; Sila-PLB . Details will be presented in a forthcoming comprehensive review ZEHPR .

### v3

@@ IV Schwarzian aspects
For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative defined for a function $f$ as Ovsienko ; Vlahakis ; ZZH ,
@@
cf. (II.6) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
@@
The form (II.10) of $S$ then implies that $P$ determines also the second solution we called $Q$.
@@
instead of $f=P/Q=1/S$ as in (IV.6).
The new cast (IV.8) then yields, once again, a SL eqn. (IV.2) whose profile is determined by the Souriau matrix,
@@
which can indeed be viewed as an equation for $S$. Somewhat surprisingly, inserting here $S$ from (II.10), the previous Sturm-Liouville equation (II.6) with the same profile $\mathcal{A}$} is recovered. These results indicate that the interchange
@@
of the Schwarzian. The difference between the two casts comes from the initial conditions. Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ involve also $S$ in (II.10).
@@ V Illustration by the Pöschl - Teller profile
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{m(m+1)}{\cosh^{2}U}\,.$ (V.1)
The Sturm-Liouville equation (II.6) is solved by Legendre functions. DM trajectories are obtained when $m$ is an integer and we get Legendre polynomials,
$P(U)=(-1)^{m}P_{m}(\tanh U)\,.$ (V.2)
The general geodesic, (III.6), combines the DM solution $P(U)$ and the non-polynomial $Q(U)$ in (III.3), shown FIG.1.
Figure 1: The solutions of the Sturm-Liouville eqn. (II.6) for Pöschl - Teller , shown for DM parameters. With initial conditions (II.19) we get Legendre polynomials, $P(U)=P_{m}(\tanh U)$. The second, independent solution, $Q(U)=Q_{m}(\tanh U)$, is a Legendre function which does not satisfy the initial condition (II.19).
Turning to BJR cf. (II.7), the Souriau matrix $S$ in (II.10) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
@@
obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, which is regular both in $I_{-}=\big\{u<0\big)\}$ and in $I_{+}=\big\{u>0\big)\}$, but diverges at $u=0$, as depicted in FIG.2a. Similarly for $m=2$ shown in FIG.2b,
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$ (V.4)
is singular where the denominator vanishes.
Figure 2: The Sturm-Liouville solutions $\bf P$ of (II.6) shown with wave numbers $\bf m=1$} and $\bf m=2$} for Pöschl - Teller in $D=1$. They yield also DM trajectories written in Brinkmann coordinates. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
$Q_{m}=P_{m}S_{m}$ in (III.3), shown in FIG.2 in green, is a second solution of the Sturm-Liouville equation (II.6). It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits VM behaviour outside the Wave zone. $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
[…]
The $P(U)$ found above provide us with DM geodesics which are consistent with (II.18). The zeros of $P$ are caustic points as it is manifest in FIG.3.
Figure 3: For a DM wave in $D=1$ transverse dimension Brinkmann trajectories focus to the caustic points determined by $P_{m}(U_{k})=0$, $k=1,\dots,m$. For $m=1$ there is one focal point, at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
Turning to symmetries, the coefficients of the globally defined Carroll generators $P$ and $Q$ in (III.1), listed in (V.6) and (V) are depicted in FIGs. 4 and 5.
@@
(a)                        (b)
Figure 4: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which generates also a trajectory] is odd. It has one zero, at $U=0$. (b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even and has two, symmetrically positioned zeros.
(a)                        (b)
Figure 5: Infinitesimal Carroll generators (V) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd and has 33 zeros.
VI Graphic representation of the symmetry generators
The symmetry-generating vectors along the trajectories shown in FIGs.4 and 5 are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory. In FIGs. 69 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of the symmetry generators $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$ in (III.1), spelt out in (V.6)-(V).
@@ VI Graphic representation of the symmetry generators
Figure 8: Translations act along the $\bf m=2$} DM trajectory, see (V). The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.8, the translation generator (with $c=1/2$) is, for $U=$-$\infty$, $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 4 and 5 confirm that, consistently with (V.6), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
@@
Boosts along the $m=2$ trajectory are shown in FIG.9. The longest vector on the left (in brown) is, consistently with (V.6) and (V) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$. Combining FIGs.4 with FIG.6 provides us with the outfolding to 3 dimensions, shown in FIG.10.
Figure 10: Translations implemented along the trajectories with wave numbers $m=1$ and $m=2$.
@@
Figure 11: Carroll boosts implemented along the trajectories with wave number $m=1$ and $m=2$.
Carroll boosts implemented along the trajectories behave as pure translations at P(U)=0P(U)~=~0. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
@@ VII Conclusion
Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.19) yields our geodesics. The general expression (III.6) which involves both DM and VM is one of our main results.
When the parameter $m$ is an integer, then the wavezone accommodates $m$ half-waves and we get DM. For $m=2\ell+1$ $P(U)$ is odd and we do get net displacement. For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.18) and confirmed by FIGs.3, 4a and 5a. The Brinkmann trajectories are focused at the zeros of $\det(P)$.
DM trajectories have zero momentum and zero non-relativistic energy DM-1 . They can be found by putting ${\bm{p}}_{0}=0$ into (III.6). Turning on the momentum, ${\bm{p}}_{0}\neq 0$, converts the trajectory to VM.
Both BJR and Brinkmann coordinates have their advantages and drawbacks.
The merit of BJR is simplicity Sou73 ; Carroll4GW . Our initial conditions (II.19) imply ${\bm{p}}=0$, and the “motion” reduces merely to a fixed point determined by the conserved boost momentum ${\bf k}={\bm{x}}_{0}$ and $v_{0}$ — the initial position (II.13)-(II.17). The dynamics is carried by $P(u)$. The price to pay is that the BJR description is valid only between the mandatory zeros of $\det(P)$, where the $S$-matrix diverges Sou73 ; LongMemory ; EZHRev , and solutions in the adjacent domains have to be fitted together EZHRev .
The advantage of the Brinkmann form (II.3)–(II.5) is that it replaces the singular Souriau matrix $S$ in (II.10) by the globally defined Sturm-Liouville solutions $P$ and $Q=PS$, (II.6) and (III.3), respectively. For zero momentum, (II.16), the non-DM solution $Q$ is turned off, leaving us with $P$ alone which plays its second, trajectory-role. The drawback is that finding those solutions may be laborious QLdPT ; DM-1 ; Sila-PLB . In conclusion, the main difficulty boils down, in both coordinate systems, to solving a Sturm-Liouville equation.
Switching to Brinkmann coordinates allowed us to extend the locally given Carrollian Killing vectors Sou73 to the whole spacetime. These vectors are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by combining the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.6), highlighting the intimate relationship of geodesics with Carroll symmetry EZHRev ; GenRelGrav .
The Brinkmann and BJR forms have also been compared by Harte Harte from a different point of view. Showing that the BJR plane wave metric is a generalisation of perturbative weak gravitational waves, he discussed memory effects in a perturbative manner.
@@
Further insight is provided by the Schwarzian framework Ovsienko ; ZZH ; Vlahakis . This approach confirms that $Q=PS$ is indeed a globally defined, 2nd solution of the Sturm-Liouville problem which does not involve the singular Souriau matrix $S$.
Our results go beyond the toy dimension $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which provides us with an analytic approximation of the physically relevant flyby case ZelPol ; Sila-PLB .
Gravitational waves and their observable effects are also formulated by means of Bondi metric Winicour which relates the gravitational emission and the mass loss. The Bondi metric is axially symmetric and asymptotically flat. In general, our plane waves are not asymptotically flat, though. However, both metrics share similar properties like having a covariantly constant null Killing vector which determines the direction of wave propagation. The transverse part of the Bondi metric is quadratic in the radial coordinate which defines the luminosity distance. In Harte , the BJR plane wave metric is used to compute similar physical properties. To clarify the differences and the similarities between the Bondi formalism with ours here requires more effort and will be reported elsewhere.
Supersymmetry MSUSY sheds further light on the Memory Effect. Further details will be presented in a comprehensive review ZEHPR .